Yasuni (04.01.2014 to 06.30.2014)

You are here

PDF versionPDF version
Year: 
2014
Quarter: 
April - June
Communications activities: 

 None

Unusual Events at the TEAM Site: 

Yasuní Site Manager Hugo Romero left PUCE in early April.
Santiago Espinosa started as new Yasuní Site Manager.

New Species: 

 None

Protocol Activities: 

Terrestrial vertebrates: The 2014 census of Array 1 and 2 was completed on May 17th (last day of camera pick up). We are finishing the processing of the images in DeskTEAM and expect to upload the TPK files of both arrays by the end of July.

Vegetation: We have finished cleaning data for the six 1-ha plots. Data will be uploaded by the end of July.

Climate: The weather station is running

fine.

 

Protocol Problems: 

Climate: We may have lost climate data for the period March 30th – May 17th. The reason is that the memory card was not retrieved properly. Hugo Romero is now checking whether the memory card has any data. 

Schedule Problems: 

Logistical Problems: 

Terrestrial vertebrates: we currently have 28 cameras working . We have sent 4 cameras to the manufacturer for repair:

Camera SN P900DD12015491: Display is not working at all

Camera SN  P900DD12115385: Battery reader does not work properly (indicates "low" with batteries fully charged)

Camera SN P900DD12115445: Camera was opened by a coati who destroyed the contacts of batteries

Camera SN P900BD07106216: We have been notified by Reconyx that this camera cannot be fixed. This camera will have to be replaced.

Climate: We need to revise our logistics for retrieving data from the climate station. We will work together with Yasuní Scientific Research Station to retrieve data and upload data on a monthly basis. 

Suggestions: 

Terrestrial vertebrates: Purchase one camera trap to replace the one that cannot be fixed